# Wilson Line and Disorder Invariants of Topological One-Dimensional Multiband Models

### v1

@@ Abstract
Topological invariants, such as the winding number, the Chern number, and the Zak phase, characterize the topological phases of bulk materials. Through the bulk-boundary correspondence, these topological phases have a one-to-one correspondence to topological edge states, which are robust to certain classes of disorder. For simple models like the Su–Schrieffer–Heeger (SSH) model, the computation of the winding number and Zak phase are straightforward, however, in multiband systems, this is no longer the case. In this work, we introduce the unwrapped Wilson line across the Brillouin zone to compute the bulk topological invariant. This method can efficiently be implemented numerically to evaluate multiband SSH-type models, including models that have a large number of distinct topological phases. This approach accurately captures all topological edge states, including those overlooked by traditional invariants, such as the winding number and Zak phase. To make a connection to experiments, we determine the sign of the topological invariant by considering a Hall-like configuration. We further introduce different classes of disorder that leave certain edge states protected, while suppressing other edge states, depending on their symmetry properties. Our approach is illustrated using different one-dimensional models, providing a robust framework for understanding topological properties in one-dimensional systems.
@@ I Introduction
Topological invariants are quantized properties of gapped Hamiltonians that remain invariant under continuous deformations. They predict the presence of symmetry-protected edge states through the bulk-boundary correspondence principle [1, 2]. Well-known topological invariants, such as the winding number, Chern number, Zak phase, and polarization, characterize bulk properties and are specific to the symmetries and dimensionality of the system [3]. Depending on the model, the class of invariant used follows from the 10-fold classification [4]. For the SSH model, the most famous of 1D topological insulators, a $\mathbb{Z}$ and $\mathbb{Z}_{2}$ invariant can be used to describe it’s topology. The winding number ($\mathbb{Z}$) directly counts zero-energy chiral symmetry-protected edge states. The Zak phase, closely related to polarization ($\mathbb{Z}_{2}$), provides an alternative approach via the Wilson loop [5]. However, these invariants have limitations. The winding number relies critically on chiral symmetry and can only count edge states pinned at zero energy, while the polarization, being a defined modulus a unit cell, cannot distinguish complex topological phases beyond 00 and $1$ [6, 7, 8]. To address these issues, we introduce a method that generalizes the Zak phase by unwrapping the Wilson line spectrum across the Brillouin zone (BZ). Unlike the Wilson loop, which is a closed path over the BZ, the Wilson line is an open-path object whose accumulated phase evolution encodes finer topological structure. By tracking the continuous phase, after unwrapping $2\pi$ discontinuities, we define a robust $\mathbb{Z}$ integer-valued topological invariant $\gamma$ applicable even when traditional invariants fail. This method is particularly useful for the multiband SSH4 model defined in section III, which lacks inversion symmetry and features edge states at non-zero energy, as it captures edge states overlooked by the winding number and polarization [9, 10, 11, 12]. In recent years, experimental studies have significantly enhanced our understanding of SSH models across multiple dimensions, with several studies reporting direct measurements of edge states in these systems [13, 14, 15]. We further examine the robustness of the edge states captured by the invariant $\gamma$ under dilute disorder, highlighting its potential relevance for future experimental exploration.
@@ II.1 Zak phase
The Zak phase is a topological invariant used to characterize one-dimensional periodic systems. It is defined as the geometric phase acquired by the electronic wavefunction as it traverses the entire Brillouin zone. It is equivalent to the Berry phase in 1D systems across a non-contractible loop [16]. The Zak phase has been shown to be related to the polarization [5] and the winding number [17]. The Zak phase is defined as follows:
@@
where $A(k)=i\braket{u(k)}{\partial_{k}}{u(k)}$ represents the Berry connection of the Bloch eigenstate |$\ket{u(k)}$. This is well-defined for single-band systems but fails in multiband cases, where the eigenstates |$\ket{u_{n}(k)}$ will have band index $n$. In a multi-band system, the Berry connection $A(k)$ generalizes to a matrix, requiring a path-ordered integral rather than a simple line integral. The Wilson loop, which we introduce next, correctly tracks the evolution of the subspace of occupied bands across the BZ.
@@ II.2 Wilson Loop Polarization
$\vartheta=\arg\prod_{i=1}^{N}\braket{u_{k_{i}}}{u_{k_{i+1}}}.$ (5)
This product defines the Wilson loop, a gauge-invariant measure computed over the Brillouin zone where |$\ket{u(k_{0})}=\ket{u(k_{N})}$. When considering multiple energy bands, the Berry phase for each segment is represented by the matrix $M_{k_{i}k_{i+1}}$:
$(M_{k_{i}k_{i+1}})_{nm}=\braket{u_{n}(k_{i})}{u_{m}(k_{i+1})},$ (6)
@@ II.3 Winding Number
[…]
@@ II.4 $\mathbb{Z}$ Wilson Line Invariant, $\gamma$
The phase of the eigenvalues $\mathcal{W}(\kappa)$ depend on the endpoint ($\kappa$) of the Wilson line:
@@
As we vary $\kappa$ from $-\pi$ to, $\pi$, we get a list of phases.
@@
The phase exhibits discontinuities with jumps of $2\pi$. We reconstruct a smooth and continuous phase function to extract the full accumulated phase across the BZ zone. To this end, we implement the use of an unwrapping process. Phase unwrapping first detects discontinuities by looking at the difference between points, then removes them by adding or subtracting multiples of $2\pi$. This process is used in applications like interferometry, where accurate phase information is necessary for precise measurements [25]. Before unwrapping, we sum the original phase over all occupied bands. The topological Wilson line invariant $\gamma$ can be calculated from the unwrapped (uw) continuous phase function $\vartheta^{\text{uw}}(\kappa)$ [26]
[…]
This expression can be computed numerically very effectively and we will use it to evaluate the topology of 1D models and confirm the bulk-boundary correspondence in all these models. We further study the robustness of the edge states to different types of disorder.
@@
1D tight-binding models play an important role in the understanding of topological properties. We will consider several generalizations of the SSH model, which we denote SSHn for generalized SSH models, where the unit cell is of size $n$ and the extended SSH model, EnSSH, of unit cell size 2, but with long-range hopping over the next $n$’th neighbor. These models are illustrated in fig. 1. We will start with the simplest topological model, the SSH model, which is equivalent to SSH2 and E0SSH in our notation.
Figure 1: Top drawing illustrates the EnSSH, which is an extended SSH model, where the extended hopping terms $t_{3}$ and $t_{4}$ extend to $n+1$ sites. In this example, $n=4$ but the unit cell has only two sites, A and B. The bottom drawing illustrates the SSHn model, where the unit cell is composed of $n$ atoms linearly coupled by $t_{i}$. In this example $n=4$ with a unit cell composed of A, B, C, and D sites.
@@ III.1 Dimer SSH
In the gauge where the second element is set to be real and positive, the corresponding Bloch eigenstates can be written as
@@
where $\varphi(k)=\arg(t_{1}+t_{2}e^{-ik})$. We will impose this gauge condition in the following analytical and numerical calculations. This gauge ensures consistency in the calculation of topological invariants by aligning the first site of the left semi-infinite chain with the A site, regardless of how many sites are present in the model. The topological invariant calculated using this gauge will give us the number of edge states on the left boundary of a right semi-infinite chain. Using either of the Bloch eigenvectors, the Zak phase index (1) becomes:
[…]
@@
[…]
Using the Wilson line in equ. (17) also yields $\gamma=\nu=\vartheta$ for this model. As previously mentioned, the SSH model has only two topological phases, 0 and 1, and its edge states can be described by both a $\mathbb{Z}_{2}$ and a $\mathbb{Z}$ topological invariant. However, this is not always the case, as is shown in the following extension of the SSH model.
@@ III.2 Extended SSH model (EnSSH)
We now illustrate the Wilson line approach by numerically computing the topological invariant of the E9SSH model using equ. (17). This model exhibits a nice series of quantized invariants from -9 to 10 as shown in fig. 3. Note that the winding number computed numerically from equ. (12) gives identical results to $\gamma$ computed using (17), which is expected. However, in the case of a multiband system, $\nu$ and $\gamma$ give distinct results (see section C). In experimental systems described by SSH-like models, the topological invariant can be identified by topological edge states using the bulk-boundary correspondence. Indeed, experimental systems are by nature finite and will generally possess two edges. For an even-length chain, the number of edge states comes in pairs (left and right edge states). Hence $\gamma=\frac{1}{2}\#$ of edges states. Numerically, we can evaluate the number of edge states from the eigenvectors and eigenvalues of the finite-size E9SSH model. For a finite system of size $N$, the onset of edge states in terms of the hopping parameters $t_{1},\cdots,t_{4}$ depends on the system size. In fig. 3 we used sizes of $N=4000$ and $N=16000$ in order to approach the infinite limit. Note that even in the SSH model, the onset of edge states is dependent on the system size, where it is given by $t_{1}/t_{2}=N/(N+2)$ [29, 30]. Due to the inversion symmetry of the EnSSH model, the number of left and right edge states is equal (for an even chain length). We find, as expected, that the total number of edge states is twice the absolute value of the Wilson line topological invariant, $\gamma$.
@@
An important property of the topological edge states, is their robustness to the presence of chiral symmetry-preserving disorder. Adding a random number $V_{\text{dis}}$ to each hopping term $t_{i}\rightarrow t_{i}+V_{\text{dis}}$ does not affect the chiral symmetry for $E=0$. Hence, zero energy edge states are preserved. For example, when looking at the topological phase $\gamma=10$ in fig. 3, adding $V_{\text{dis}}$ doesn’t affect the $202\gamma=$ 20 edge states at $E=0$. This behaviour is similar to many SSH-like models [31]. However, the disorder cannot be too large, otherwise the bulk gap closes.
The number of edge states only gives the absolute value of $\gamma$ (divided by two). It would be interesting to find a method, which could measure the sign change of $\gamma$ as shown in fig. 3. In two-dimensional systems, the toplogical Chern number, can be measured by the Hall conductance in the quantum Hall regime. In one-dimensional models, there is no Hall conductance, but the unit cell of the extended SSH models has 2 atoms, so it is possible to measure the transverse voltage along the atom unit cell as illustrated in the fig. 4b. Since the sign of $\gamma$ is related to the relative phase of the A and B atoms of the unit cell, we can extract the sign of $\gamma$ in an equivalent four-terminal Hall configuration. To illustrate this point, we computed the Hall resistance (transverse resistance) using a finite-sized E3SSH model where the voltage contacts are weakly coupled to the chain. We have contact probes, $V_{S}$ and $V_{D}$, which are the source and drain contacts, and $V_{1}$ and $V_{2}$ for the transverse Hall contacts. We used the standard four-terminal Fisher-Lee expression to evaluate the Hall resistance [32, 33]. For a finite chain of length of $N$ given by Hamiltonian $H_{N}$, the four-terminal Hall resistance configuration in this formalism is given by $R_{H}=R_{2,S}-R_{3,S}$ ($S$ the source, 1,2 the voltage contacts, and $D$ the drain), where
@@
$T_{ij}=\text{Tr}(\Gamma_{i}G\Gamma_{j}G^{\dagger})$, $\Gamma_{i}=\Sigma_{i}-\Sigma_{i}^{\dagger}$, and $G=(E-H_{N}-\sum_{i}\Sigma_{i})^{-1}$. The self-energies are zero matrices of the size of $H_{N}$ with the non-zero diagonal element $\Sigma_{i}=|t_{L}|^{2}e^{ik}$ at the position of the leads ($i=S,1,2,\text{ or }D$), where $E=2\cos(k)$. We assumed semi-infinite leads of hopping element unity and coupled to the EnSSH chain by $t_{L}$. Using wide band leads doesn’t change the results. In Fig. 4c we show the Hall resistance as a function of energy together with the Wilson line invariant computed using equ. (17). The Hall resistance is computed for a finite chain length of N=800, while $\gamma$ is obtained for the periodic system. Only in the limit $N\rightarrow\infty$ do we expect the position of the different phases to match. However, even with $N=800$ we see that the agreement is already very good and the sign of the Hall resistance is negative when $\gamma$ is negative and vice versa. Note that we do not expect a quantization of the Hall resistance, unlike what is observed in two-dimensional topological insulators. Nevertheless, this Hall configuration would allow the determination of the sign change of $\gamma$. Hence, combining this Hall configuration with a measurement of the number of edge states, would fully characterize the topology of an extended SSH model.

### v2

@@ Abstract
For simple models like the Su–Schrieffer–Heeger (SSH) model, the characterization of its topology via the winding number and Zak phase is straightforward; however, in multiband systems, this is no longer the case. In this work, we introduce the unwrapped Wilson line across the Brillouin zone to compute the bulk topological invariant. This method can be efficiently implemented numerically to evaluate multiband SSH-type models, including models that have a large number of distinct topological phases. This approach accurately captures all topological edge states, including those overlooked by traditional invariants, such as the winding number and Zak phase. To make a connection with the experiments, we determine the sign of the topological invariant by considering a Hall-like configuration. We further introduce different classes of disorder that leave certain edge states protected, while suppressing other edge states, depending on their symmetry properties. Our approach is illustrated using different one-dimensional models, providing a robust framework for understanding topological properties in one-dimensional systems.
@@ I Introduction
Topological invariants are quantized properties of gapped Hamiltonians that remain invariant under continuous deformations. They predict the presence of symmetry-protected edge states through the bulk-boundary correspondence principle [1, 2]. Well-known topological invariants, such as the winding number, Chern number, Zak phase, and polarization, characterize bulk properties and are specific to the symmetries and dimensionality of the system [3]. Depending on the model, the class of invariant used follows from the 10-fold classification [4]. For the SSH model, the most famous of 1D topological insulators, a $\mathbb{Z}$ and $\mathbb{Z}_{2}$ invariant can be used to describe it’s topology. The winding number ($\mathbb{Z}$) directly counts zero-energy chiral symmetry-protected edge states. The Zak phase, closely related to polarization ($\mathbb{Z}_{2}$), provides an alternative approach via the Wilson loop [5]. However, these invariants, have limitations; the winding number relies critically on chiral symmetry and can only count edge states pinned at zero energy, while the polarization, being a defined modulus a unit cell, cannot distinguish complex topological phases beyond 0 and $1$ [6, 7, 8]. To address these issues, we introduce a method that generalizes the Zak phase by unwrapping the Wilson line spectrum across the Brillouin zone (BZ). Unlike the Wilson loop, which is a closed path over the BZ, the Wilson line is an open-path object whose accumulated phase evolution encodes finer topological structure. By tracking the continuous phase, after unwrapping $2\pi$ discontinuities, we define a robust $\mathbb{Z}$ integer-valued topological invariant $\gamma$ applicable even when traditional invariants fail.
This method is particularly useful for the multiband SSH4 model defined in section III, which lacks inversion symmetry and features edge states at non-zero energy which are overlooked by the winding number and polarization [9, 10, 11, 12]. Models with inversion symmetry have been discussed in detail in [13], as well as non-zero energy edge states. Although these non-zero edge states within the band subgap of the SSH4 do not have a gap closing to indicate a topological transition, works such as [2] suggest that a change in the symmetry responsible for protecting said states can also signify a topological phase transition.
In recent years, experimental studies have significantly improved our understanding of SSH models across multiple dimensions, with several studies reporting direct measurements of edge states in these systems [14, 15, 16]. We further examine the robustness of the edge states captured by the invariant $\gamma$ under dilute disorder, highlighting its potential relevance for future experimental exploration.
@@ II.1 Zak phase
The Zak phase is a topological invariant used to characterize one-dimensional (1D) periodic systems. It is defined as the geometric phase acquired by the electronic wavefunction as it traverses the entire Brillouin zone. It is equivalent to the Berry phase in 1D systems across a non-contractible loop [17]. The Zak phase has been shown to be related to polarization [5] and winding number [18]. The Zak phase is defined as follows:
$Z=\int_{-\pi}^{\pi}A(k)dk,$ (1)
@@ II.2 Wilson Loop Polarization
$\vartheta=\arg\prod_{i=1}^{N}\braket{u_{k_{i}}|u_{k_{i+1}}}.$ (5)
[…]
$(M_{k_{i}k_{i+1}})_{nm}=\braket{u_{n}(k_{i})|u_{m}(k_{i+1})},$ (6)
@@ II.3 Winding Number
$\nu=\frac{i}{2\pi}\int_{-\pi}^{\pi}\text{Tr}\left(Q^{-1}(k)\frac{\partial Q(k)}{\partial k}\right)\,dk.$ (12)
@@ II.4 $\mathbb{Z}$ Wilson Line Invariant, $\gamma$
The phase of the eigenvalues $\mathcal{W}(\kappa)$ depends on the endpoint ($\kappa$) of the Wilson line:
@@
As we vary $\kappa$ from $\pi$ to $\pi$, we get a list of phases.
@@
The phase exhibits discontinuities with jumps of $2\pi$. We reconstruct a smooth and continuous phase function to extract the full accumulated phase across the BZ zone. To this end, we implement the use of an unwrapping process. Phase unwrapping first detects discontinuities by looking at the difference between points, then removes them by adding or subtracting multiples of $2\pi$. This process is used in applications such as interferometry, where accurate phase information is necessary for precise measurements [26]. Before unwrapping, we sum the original phase over all occupied bands. The topological Wilson line invariant $\gamma$ can be calculated from the unwrapped (uw) continuous phase function $\vartheta^{\text{uw}}(\kappa)$ [27]
$\gamma=\frac{1}{\pi}\int_{-\pi}^{\pi}\frac{\partial}{\partial\kappa}\vartheta^{\text{uw}}(\kappa)\,d\kappa=(\vartheta^{\text{uw}}(\pi)-\vartheta^{\text{uw}}(-\pi)).$ (17)
The unwrapping process results in $\gamma$ being some integer of $2\pi$ beyond the Wilson Loop Zak phase. Therefore, whenever the Zak phase is quantized, that is, when chiral or inversion symmetry is present, $\gamma$ will also be quantized. It is important to note that $\gamma$ is gauge dependent. We choose a gauge such that the first or last element of $u(k)$ is real. If taking a gauge where the last element is real, consistency is ensured by aligning the first site of the right semi-infinite chain with a $A$ site in real space, regardless of how many sites are present in the model. This gauge is equivalent to setting the virtual site $D_{i}=0$ where the site index $i=0$, so that no electron can hop to the left of $A_{1}$. The topological invariant calculated using this gauge will give us the number of edge states on the left boundary of a right semi-infinite chain. If choosing the first element of the eigenstate to be real, the same logic applies, and $\gamma$ counts the number of right edge states of a left semi-infinite chain. The Wilson line invariant can be computed numerically very effectively, and we will use it to evaluate the topology of 1D models and confirm the bulk-boundary correspondence. We further study the robustness of the topological edge states to different types of disorder.
@@
1D tight-binding models play an important role in understanding topological properties. We will consider several generalizations of the SSH model belonging to the BDI symmetry class. The SSHn model generalizes the SSH chain to a unit cell of size $n$, while the extended SSH model , EnSSH, has a unit cell of two sites (A and B) but allows long-range hopping over a distance of $n$ unit cells, always between A and B sites to preserve chiral symmetry.These models are illustrated in fig. 1. We will start with the simplest topological model, the SSH model, which is equivalent to SSH2 and E0SSH in our notation.
Figure 1: Top drawing illustrates the EnSSH, which is an extended SSH model with two sites, A and B, per unit cell. In this example, the E2SSH is plotted. The bottom drawing illustrates the SSHn model, where the unit cell is composed of $n$ atoms linearly coupled by $t_{i}$. In this example, the SSH4 has a unit cell composed of A, B, C, and D sites.
@@ III.1 Dimer SSH
In the gauge, where the second element is set to be real and positive, the corresponding Bloch eigenstates can be written as
@@
where $\varphi(k)=\arg(t_{1}+t_{2}e^{-ik})$. We will impose this gauge condition in the following analytical and numerical calculations. This gauge ensures consistency in the calculation of topological invariants by aligning the first site of the left semi-infinite chain with the $A_{1}$ site, regardless of how many sites are present in the model. This gauge is equivalent to setting the virtual site $D_{0}=0$, so that no electron can hop to the left of $A_{1}$. The topological invariant calculated using this gauge will give us the number of edge states on the left boundary of a right semi-infinite chain. Using either of the Bloch eigenvectors, the Zak phase index (1) becomes:
$\vartheta=\frac{Z}{\pi}$ $=-\frac{1}{2\pi}\int_{-\pi}^{\pi}\frac{\partial\varphi(k)}{\partial_{k}}dk$
$=-\frac{1}{2\pi i}\int_{-\pi}^{\pi}\frac{\partial}{\partial_{k}}\mathrm{ln}(t_{1}+t_{2}e^{-ik})dk$ (20)
@@
$\nu=\frac{i}{2\pi}\int_{-\pi}^{\pi}\frac{1}{t_{1}+t_{2}e^{-ik}}\frac{\partial(t_{1}+t_{2}e^{-ik})}{\partial k}dk=\vartheta.$ (22)
Using the Wilson line in equ. (17) also yields $\gamma=\nu=\vartheta$ for this model. As mentioned above, the SSH model has only two topological phases, 0 and 1, and its edge states can be described by both a $\mathbb{Z}_{2}$ and a $\mathbb{Z}$ topological invariant. However, this is not always the case, as is shown in the following extension of the SSH model.
@@ III.2 Extended SSH model (EnSSH)
We now illustrate the Wilson line approach by numerically computing the topological invariant of the E9SSH model using equ. (17). Here E9SSH means $A_{i}\!\to\!B_{i+9}$ and $B_{i}\!\to\!A_{i+10}$, and all computations use $q(k)$ from equ. (24). This model exhibits a nice series of quantized invariants from -9 to 10 as shown in fig. 3. Note that the winding number computed numerically from equ. (12) gives identical results to $\gamma$ computed using (17), which is expected. However, in the case of a multiband system, $\nu$ and $\gamma$ give distinct results (see section C). In experimental systems described by SSH-like models, the topological invariant can be identified by topological edge states using the bulk-boundary correspondence. Indeed, many experimental systems are finite and possess two edges. For an even-length chain, the number of edge states comes in pairs (left and right edge states). Hence $\gamma=\frac{1}{2}\#$ of edge states. Numerically, we can evaluate the number of edge states from the eigenvectors and eigenvalues of the finite-size E9SSH model. For a finite system of size $N$, the onset of edge states in terms of the hopping parameters $t_{1},\cdots,t_{4}$ depends on the size of the system. In fig. 3 we used sizes of $N=4000$ and $N=16000$ in order to approach the infinite limit. Note that even in the SSH model, the onset of edge states depends on the size of the system, where it is given by $t_{1}/t_{2}=N/(N+2)$ [30, 31]. Due to the inversion symmetry of the EnSSH model, the number of left- and right edge states is equal (for an even chain length). We find, as expected, that the total number of edge states is twice the absolute value of the Wilson line topological invariant, $\gamma$.
@@
An important property of the topological edge states, is their robustness to the presence of chiral symmetry-preserving disorder. Adding a random number $V_{\text{dis}}$ to each hopping term $t_{i}\rightarrow t_{i}+V_{\text{dis}}$ does not affect the chiral symmetry for $E=0$. Hence, zero energy edge states are preserved. For example, when looking at the topological phase $\gamma=10$ in fig. 3, adding $V_{\text{dis}}$ doesn’t affect the $2\gamma=20$ edge states at $E=0$. This behavior is similar to many SSH-like models [32]. However, the disorder cannot be too large, otherwise, the bulk gap closes.
The number of edge states only gives the absolute value of $\gamma$ (divided by two). It would be interesting to find a method, which could measure the sign change of $\gamma$ as shown in fig. 3. In two-dimensional systems, the toplogical Chern number, can be measured by the Hall conductance in the quantum Hall regime. In 1D models, there is no Hall conductance, but the unit cell of the extended SSH models has 2 atoms, so it is possible to measure the transverse voltage along the atom unit cell as illustrated in the fig. 4b. Since the sign of $\gamma$ is related to the relative phase of the A and B atoms of the unit cell [33], we can extract the sign of $\gamma$ in an equivalent four-terminal Hall configuration. To illustrate this point, we computed the Hall resistance (transverse resistance) using a finite-sized E3SSH model where the voltage contacts are weakly coupled to the chain. We have contact probes, $V_{S}$ and $V_{D}$, which are the source and drain contacts, and $V_{1}$ and $V_{2}$ for the transverse Hall contacts. We used the standard four-terminal Fisher-Lee expression to evaluate the Hall resistance [34, 35]. For a finite chain of length $N$ given by the Hamiltonian $H_{N}$, the four-terminal Hall resistance configuration in this formalism is given by $R_{H}=R_{2,S}-R_{1,S}$ ($S$ the source, 1,2 the voltage contacts, and $D$ the drain), where
@@
$T_{ij}=\text{Tr}(\Gamma_{i}G\Gamma_{j}G^{\dagger})$, $\Gamma_{i}=\Sigma_{i}-\Sigma_{i}^{\dagger}$, and $G=(E-H_{N}-\sum_{i}\Sigma_{i})^{-1}$. The self-energies are zero matrices of the size of $H_{N}$ with the non-zero diagonal element $\Sigma_{i}=|t_{L}|^{2}e^{ik}$ at the position of the leads ($i=S,1,2,\text{ or }D$), where $E=2\cos(k)$. We assumed semi-infinite leads of hopping element unity and coupled to the EnSSH chain by $t_{L}$. Using wide band leads doesn’t change the results. In Fig. 4c we show the Hall resistance as a function of energy together with the Wilson line invariant computed using equ. (17). The Hall resistance is computed for a finite chain length of N=800, while $\gamma$ is obtained for the periodic system. Only in the limit $N\rightarrow\infty$ do we expect the positions of the different phases to coincide. However, even with $N=800$ we see that the agreement is already very good and the sign of the Hall resistance is negative when $\gamma$ is negative and vice versa. Note that we do not expect a quantization of the Hall resistance, unlike what is observed in two-dimensional topological insulators. Nevertheless, this Hall configuration would allow for the determination of the sign change of $\gamma$. Hence, combining this Hall configuration with a measurement of the number of edge states would fully characterize the topology of an extended SSH model.